# Failures to Surface Harmful Contents in Video Large Language Models

### v1

@@ Abstract
I Introduction
Video Large Language Models (VideoLLMs) have recently become state-of-the-art engines for video understanding [1, 2, 3]. They distill high-level semantics from diverse footage, such as classroom lectures, tutorials, news segments, sports highlights, entertainment shows, surveillance clips, and more, then generate concise summaries or detailed textual interpretations. By condensing lengthy footage into concise textual summaries, VideoLLMs condense lengthy footage into concise textual summaries, enabling viewers to skim the video casually while relying on the generated text to grasp its main ideas. This new way of video consumption markedly improves accessibility and eases cognitive load, making VideoLLMs indispensable for students, professionals, content moderators, and general users [4].
@@ I Introduction
To investigate this issue, we dissect VideoLLMs’ processing pipeline and identify three structural flaws that give rise to this semantic blind spot. First, VideoLLMs typically adopt sparse uniform frame sampling to keep computation tractable. This leaves large portions of the video unexamined, allowing attackers to insert harmful content in unsampled intervals without detection [5]. Second, the retained frames often undergo aggressive spatial downsampling to trim visual tokens [6], which leads to the loss of fine-grained information from small regions, such as a small corner patch. Third, the cross-modal decoder downplays visual evidence: linguistic priors dominate the attention budget, so cues that do survive tokenization may still be ignored at generation time [7]. Combined, these three structural flaws, temporal sparse sampling, spatial downsampling, and modality fusion imbalance, collectively account for VideoLLMs’ consistent omission of harmful content. Motivated by these findings, we craft three zero-query, black-box omission attacks, each exploiting one or more of these flaws:
@@
To quantify the severity of the omission vulnerability, we comprehensively test the three proposed attacks against five representative VideoLLMs, LLaVA-Video-7B-Qwen2 [8], LLaVA-NeXT-Video-7B-DPO [9], LLaVA-NeXT-Video-32B-Qwen [9], VideoLLaMA2 [10], and ShareGPT4Video [11], using test clips that embed three types of harmful content: violence, crime, and pornography. Using the unified metric of Harmful Omission Rate (HOR), the percentage of harmful clips that pass unmentioned, we find that, with hyperparameters ensuring the harmful content remains clearly visible and semantically recognizable to human viewers, the average HORs remain strikingly high: 99%, 91%, 100% for violence, crime and pornography content under FRA; 98%, 87%, 76% under PPA; and 93%, 82%, 93% under TOA. This phenomenon reflects the fact that most injected frames evade sampling due to temporal sparsity, and harmful content in corners is largely discarded during spatial downsampling. Even when some visual tokens survive, they are often suppressed by unbalanced modality fusion, a failure that also occurs in TOA, where transparent yet clearly visible harmful content is added to every frame. These findings reveal a fundamental weakness in state-of-the-art VideoLLMs and underscore the need for denser temporal sampling, finer spatial token retention, and a more balanced cross-modal fusion to achieve reliable safety against harmful content. Our code is provided in the supplement and will be made publicly available upon acceptance.
@@
  • We comprehensive test these three attacks against five representative VideoLLMs with three types of harmful videos, and the results highlight the severity of the vulnerability and underscore the urgent need for VideoLLMs to advance their fundamental design.
II Background
@@
Pretrained on large-scale image-text datasets, the visual encoder, such as SigLIP [12], BLIP-2 [13] and EVA-CLIP [14], extracts visual embeddings from a uniformly sampled subset of frames from the input video. These embeddings are then mapped by the projector into the same embedding space as the language input tokens. Projectors are typically implemented using Multi-Layer Perceptrons, cross-attention modules [15], or Q-Formers [13]. The LLM serves as the core reasoning engine. It receives a concatenated sequence of projected visual tokens and text tokens, and produces the textual output. Most VideoLLMs use instruction-tuned models such as LLaMA [16], Vicuna [17] and Qwen [18] as their backbone. This unified design enables VideoLLMs to jointly process visual and textual inputs, supporting diverse video understanding tasks including video captioning and question answering.
@@ III Related Work
Video Large Language Models. Early progress in multimodal LLMs (MLLMs), such as Flamingo [19] and BLIP-2 [13], shows that pairing LLMs with visual encoders can excel at image-based tasks such as captioning and visual question answering [20, 16]. The same demand for rich, multimodal reasoning now extends to the temporal domain, spurring the rise of VideoLLMs that aim to interpret dynamic, semantically complex video content. To cope with the higher spatio-temporal complexity, modern VideoLLMs extract spatial-temporal features from video frames, align them with language embeddings, and feed the combined embeddings into a pretrained LLM [6]. Since processing full videos is prohibitively expensive on GPU memory and compute limits, most existing systems (e.g., Video-LLaMA2 [10], InternVL2.5 [21] and NVILA [22]) resort to sparse uniform sampling. ViLaMP [23] suggests that sampling 16 frames for short videos and 32 for long ones provides a good trade-off between efficiency and performance. However, fixed low sampling rate regardless of video length results in uneven temporal spacing (the interval grows with video duration), and causes critical segments to be skipped, leaving large portions of the video unexamined. A few models, such as ShareGPT4Video [11], VideoAgent [24] and AKS [25], further apply key frame selection, yet the final number of frames remains small.
@@
Moreover, token-level compression techniques are also widely used. Typically, the visual tokens are downsampled using a $2\times 2$ bilinear interpolation in LLaVA-OneVision [6] and VideoLLaMA3 [26], or average pooling in LLaVA-Video [8]. Some other models reduce visual tokens through various compression strategies. For instance, LLaMA-VID [27] fixes the number of tokens per frame to two, NVILA [22] scales up spatial and temporal resolution before pooling, and Chat-UniVi [28] performs k-nearest-neighbor based clustering to reduce redundancy. However, such token compression strategies result in extremely limited information per frame, leading to the loss of fine-grained visual details. More details of existing mainstream VideoLLMs are provided in Appendix.
@@
MLLM Safety. The safety of MLLMs has become a pressing concern for surveillance, content moderation, and educational applications, where models must reliably recognize and react to harmful material such as violence, nudity, or abuse. Recent studies [7] reveal that image MLLMs often underutilize visual features during decoding. Even when meaningful signals are extracted by the visual encoder, the fusion and decoding stages tend to favor linguistic priors. Despite growing attention to similar safety risks in image MLLMs [29, 30] and generative video models [31, 32], safety vulnerabilities in VideoLLMs remain largely underexplored. To bridge this gap, we systematically characterize omission failures in VideoLLMs and introduce corresponding attacks that expose fundamental flaws in their design.
IV Analyses
@@
Flaw 2: Token Under-Sampling. Modern VideoLLMs inherit the input token limit of their host LLMs. For example, GPT-4 allows at most 8,192 tokens per input [33]. Since this token budget is shared between visual tokens and textual tokens, VideoLLMs must compresses the tokens of each frame to meet the token budget limit. Formally, given a video of $N$ sampled frames and per-frame token number $P$, the total token number should satisfy:
@@
where $B$ denotes the LLM’s input token limit. Therefore, the token number is reduced to $P^{\prime}$ ($P^{\prime}<P$) through token compression. Many recent works adopt simple downsampling techniques, such as bilinear interpolation [26, 6] or average pooling [8], which aggregates visual tokens on a 2D spatial token grid. For example, an original $14\times 14$ token grid is downsampled to $7\times 7$, retaining only 25% of the spatial tokens.
@@
Flaw 3: Modality Fusion Imbalance. After projection into the language model’s embedding space, visual tokens are often underutilized during decoding. As a result, the LLM tends to prioritize textual information while downplaying or even ignoring signals from the visual encoder, preventing harmful cues from being reflected in the final response. Prior studies [7] show that the standalone visual encoder surpasses the fully fused image-text model on vision-centric benchmarks, underscoring a structural imbalance where visual information loses influence after fusion and barely shapes the final representation. The problem persists in existing VideoLLMs, which reuse image-based encoders to generate visual tokens. Even when these encoders flag harmful content in the sampled frames, their signals are diminished during decoding, hindering the model from faithfully reporting it.
To validate this, we conduct a comparative experiment using LLaVA-Video-7B-Qwen2 and its underlying visual encoder, SigLIP. Following [7], we examine the effectiveness of the visual encoder through a visual probing strategy. Specifically, we construct evaluation videos by inserting harmful content into benign source videos (details in the next section) and randomly sample 100 benign and 100 harmful examples for each of three categories: violence, crime, and pornography. For each video, we examine the binary classification accuracy of both the standalone visual encoder and the full VideoLLM under identical inputs. Figure 3 reports the proportion of correctly identified videos for each category. For benign videos, both the visual encoder and the full VideoLLM achieve comparably high detection rates. However, for the three harmful categories, the full VideoLLM exhibits a significant performance drop compared to the visual encoder. This discrepancy provides concrete empirical support for this flaw, confirming that modality fusion does suppress visual signals even when they are preserved at the visual encoder level.
@@ V Attack Approaches
V-A Threat Model
We assume a strict zero-query black-box setting: the adversary has no knowledge of the targeting VideoLLM’s internals, such as architecture, weights, training data, temporal sampling rate, token ompression strategy, or modality fusion scheme, and cannot repeatedly query the model for optimization. The only prior is the knowledge of the three architectural flaws identified previously. The adversary may insert self-sourced harmful clips, but these must remain visible to a human i.e., not single-frame flashes or imperceptible perturbations, so that any detection failure reflects a true omission. This zero-query, training-free attacking setup enables efficient, real-time deployment without per-video adaptation.
V-B Three Attacks
@@
Frame-Replacement Attack (FRA $\rightarrow$ Flaw 1, 3). This attack replaces a segment of the original video with a harmful video clip at a randomly chosen position. Specifically, we select a random insertion point and overwrite the subsequent $t_{r}$ seconds ($t_{r}>1$) with a preselected harmful video clip. Since VideoLLMs employ the sparse uniform sampling, a short replacement window $t_{r}$ is seldom sampled. For instance, in a 2-minute video at 30 FPS, i.e., 3600 frames, taking only 16 evenly spaced frames gives a stride of 8 seconds (240 frames). A 4-second harmful clip can therefore fit entirely between two sampled frames, leaving the model with no evidence of it, even though human viewers see this harmful segment clearly.
Picture-in-Picture Attack (PPA $\rightarrow$ Flaw 2, 3). We particularly embed a harmful clip in a fixed Picture-in-Picture (PiP) region within each frame, e.g., the bottom-right corner, while the rest of the frame remains unchanged. The PiP region occupies $\eta H\times\eta W$ pixels of each frame, where $\eta\in(0,1)$ is chosen not too small to ensure the malicious content is visible to humans. Because VideoLLMs compress the tokens of each frame to fit the token budget, small peripheral regions are often discarded, failing to influence the model’s output despite being clearly visible.
@@
VI Experiments
VI-A Experimental Setup
Video Samples. We randomly sample 200 original videos from the LLaVA-Video-178K dataset [8], a widely used benchmark for evaluating VideoLLMs. For harmful clips, we focus on three representative categories: violence, crime, and pornography. These categories are commonly encountered in safety-critical scenarios and easily recognizable to humans. Harmful videos are collected from public datasets (RLVS [34], XD-Violence [35], Pornography dataset [36]) and online platforms including YouTube and Pornhub. For each category, we randomly select 10 harmful clips for the three attacks. For each attack, we randomly pair every original video with one harmful video from the corresponding category.
@@ VI-A Experimental Setup
VI-B Experimental Results
Attack Effectiveness. Table I demonstrates the effectiveness of the three zero-query black-box attacks. For each attack, we fix a representation hyperparameter settings that are chosen to keep the harmful content clearly visible to humans while inducing substantial omissions by VideoLLMs. Attackers can readily tune these parameters to suit their own objectives.
@@ VI-B Experimental Results
Visualizations. Figure 5 illustrates representative video examples for all three attacks and harmful content categories. In every case, the injected clip is clearly visible to humans, yet every evaluated model fails to mention it. These omissions reveal the vulnerability of current VideoLLMs to harmful content injection, driven by their fundamental weakness.
@@
VI-C Detailed Prompt Analysis on Failure Cases
In our experiments, we use HOR as the primary metric to assess whether the model could detect harmful content. However, further analysis reveals that even in cases where attacks are deemed unsuccessful (i.e., the VideoLLM provides a positive response to our prompt), the model still struggle when presented with more detailed question regarding the harmful content. In particular, we query the model with the following prompt: “Please check if the video contains crime. If yes, specify the time range (start and end) in minutes and seconds, spatial location (one of center, corner, full frame), and briefly describe the scene.”
Table II presents several representative examples. We find that the model often fails to provide accurate responses regarding the temporal occurrence, spatial location, or the nature of the harmful content. Specifically, in all cases, the VideoLLM is unable to correctly identify the time range where the harmful content occurs. More surprisingly, many time ranges output by the model exceed the actual duration of the video, indicating that the model cannot temporally locate specific segments. This limitation is understandable, as the model operates on sparsely sampled frames and therefore lacks coherent access to the temporal sequence. For spatial localization, the model performs relatively well when identifying harmful content located in the corner. However, it struggles to distinguish between full-frame and center regions. This limitation can be attributed to excessive token compression applied during visual encoding, which tends to blur fine-grained spatial information and reduce the model’s ability to preserve positional distinctions. As a result, the model’s spatial localization ability remains coarse and lacks precision. Regarding scene-level descriptions of harmful content, the model frequently produces responses that do not match the inserted harmful content at all. In most cases, the description corresponds to benign elements from the original video (which, notably, the model does not consider harmful when provided alone), or it refers to hallucinates content that never appears in the video, such as in Case 3 of PPA.
TABLE II: Examples of failure cases under more detailed prompt. The model correctly detects the presence of harmfulness but fails to localize or describe it accurately. R: Real information. M: Model output. Red denotes that the model output is inaccurate.

### v2

@@ Abstract
Introduction
Video Large Language Models (VideoLLMs) have recently become state-of-the-art engines for video understanding (zhao2023learning; tang2025video; weng2024longvlm). They distill high-level semantics from diverse footage, such as classroom lectures, tutorials, news segments, sports highlights, entertainment shows, surveillance clips, and more, then generate concise summaries or detailed textual interpretations. By condensing lengthy footage into concise textual summaries, VideoLLMs enable viewers to skim the video casually while relying on the generated text to grasp its main ideas. This new way of video consumption markedly improves accessibility and eases cognitive load, making VideoLLMs indispensable for students, professionals, content moderators, and general users (qian2024streaming).
@@ Introduction
To investigate this issue, we dissect VideoLLMs’ processing pipeline and identify three structural flaws that give rise to this semantic blind spot. First, VideoLLMs typically adopt sparse uniform frame sampling to keep computation tractable. This leaves large portions of the video unexamined, allowing attackers to insert harmful content in unsampled intervals without detection (li2025improving). Second, the retained frames often undergo aggressive spatial downsampling to trim visual tokens (li2024llava-onevision), which leads to the loss of fine-grained information from small regions, such as a small corner patch. Third, the cross-modal decoder downplays visual evidence: linguistic priors dominate the attention budget, so cues that do survive tokenization may still be ignored at generation time (fu2025hidden). Combined, these three structural flaws, temporal sparse sampling, spatial downsampling, and modality fusion imbalance, collectively account for VideoLLMs’ consistent omission of harmful content. Motivated by these findings, we craft three zero-query, black-box omission attacks, each exploiting one or more of these flaws:
@@
To quantify the severity of the omission vulnerability, we comprehensively test the three proposed attacks against five representative VideoLLMs, LLaVA-Video-7B-Qwen2 (zhang2024LLaVA-Video), LLaVA-NeXT-Video-7B-DPO (zhang2024llavanextvideo), LLaVA-NeXT-Video-32B-Qwen (zhang2024llavanextvideo), VideoLLaMA2 (cheng2024videollama2), and ShareGPT4Video (chen2024sharegpt4video), using test clips that embed three types of harmful content: violence, crime, and pornography. Using the unified metric of Harmfulness Omission Rate (HOR), the percentage of harmful clips that pass unmentioned, we find that, with hyperparameters ensuring the harmful content remains clearly visible and semantically recognizable to human viewers, the average HORs remain strikingly high: 99%, 91%, 100% for violence, crime and pornography content under FRA; 98%, 87%, 76% under PPA; and 93%, 82%, 93% under TOA. This phenomenon reflects the fact that most injected frames evade sampling due to temporal sparsity, and harmful content in corners is largely discarded during spatial downsampling. Even when some visual tokens survive, they are often suppressed by unbalanced modality fusion, a failure that also occurs in TOA, where transparent yet clearly visible harmful content is added to every frame. These findings reveal a fundamental weakness in state-of-the-art VideoLLMs and underscore the need for denser temporal sampling, finer spatial token retention, and a more balanced cross-modal fusion to achieve reliable safety against harmful content. Our code is available at https://github.com/yuxincao22/VideoLLM-Failures.
@@
  • We comprehensive test these three attacks against five representative VideoLLMs with three types of harmful videos, and the results highlight the severity of the vulnerability and underscore the urgent need for VideoLLMs to advance their design.
Background
@@
Pretrained on large-scale image-text datasets, the visual encoder, such as SigLIP (zhai2023siglip), BLIP-2 (li2023blip) and EVA-CLIP (fang2023eva), extracts visual embeddings from a uniformly sampled subset of frames from the input video. These embeddings are then mapped by the projector into the same embedding space as the language input tokens. Projectors are typically implemented using Multi-Layer Perceptrons, cross-attention modules (vaswani2017attention), or Q-Formers (li2023blip). The LLM serves as the core reasoning engine. It receives a concatenated sequence of projected visual tokens and text tokens, and produces the textual output. Most VideoLLMs use instruction-tuned models such as LLaMA (touvron2023llama), Vicuna (chiang2023vicuna) and Qwen (bai2023qwen) as their backbone. This unified design enables VideoLLMs to jointly process visual and textual inputs, supporting diverse video understanding tasks including video captioning and question answering.
@@
Video Large Language Models. Early progress in multimodal LLMs (MLLMs), such as Flamingo (alayrac2022flamingo) and BLIP-2 (li2023blip), shows that pairing LLMs with visual encoders can excel at image-based tasks such as captioning and visual question answering (radford2018improving; touvron2023llama). The same demand for rich, multimodal reasoning now extends to the temporal domain, spurring the rise of VideoLLMs that aim to interpret dynamic, semantically complex video content. To cope with the higher spatio-temporal complexity, modern VideoLLMs extract spatial-temporal features from video frames, align them with language embeddings, and feed the combined embeddings into a pretrained LLM (li2024llava-onevision). Since processing full videos is prohibitively expensive on GPU memory and compute limits, most existing systems (e.g., Video-LLaMA2 (cheng2024videollama2), InternVL2.5 (chen2024InternVL2.5) and NVILA (liu2025nvila)) resort to sparse uniform sampling. ViLaMP (cheng2025vilamp) suggests that sampling 16 frames for short videos and 32 for long ones provides a good trade-off between efficiency and performance. However, fixed low sampling rate regardless of video length results in uneven temporal spacing (the interval grows with video duration), and causes critical segments to be skipped, leaving large portions of the video unexamined. A few models, such as ShareGPT4Video (chen2024sharegpt4video), VideoAgent (fan2024videoagent) and AKS (tang2025adaptive), further apply key frame selection, yet the final number of frames remains small.
@@
Moreover, token-level compression techniques are also widely used. Typically, the visual tokens are downsampled using a $2\times 2$ bilinear interpolation in LLaVA-OneVision (li2024llava-onevision) and VideoLLaMA3 (zhang2025videollama3), or average pooling in LLaVA-Video (zhang2024LLaVA-Video). Some other models reduce visual tokens through various compression strategies. For instance, LLaMA-VID (li2024llama-vid) fixes the number of tokens per frame to two, NVILA (liu2025nvila) scales up spatial and temporal resolution before pooling, and Chat-UniVi (jin2024chat-univi) performs k-nearest-neighbor based clustering to reduce redundancy. However, such token compression strategies result in extremely limited information per frame, leading to the loss of fine-grained visual details. More details of existing mainstream VideoLLMs are provided in Appendix.
@@
MLLM Safety. The safety of MLLMs has become a pressing concern for surveillance, content moderation, and educational applications, where models must reliably recognize and react to harmful material such as violence, nudity, or abuse. Recent studies (fu2025hidden) reveal that image MLLMs often underutilize visual features during decoding. Even when meaningful signals are extracted by the visual encoder, the fusion and decoding stages tend to favor linguistic priors. Despite growing attention to similar safety risks in image MLLMs (liu2024safety_ijcai; ying2024safebench) and generative video models (wang2024gpt4video; chen2024safewatch), safety vulnerabilities in VideoLLMs remain largely underexplored. To bridge this gap, we systematically characterize omission failures in VideoLLMs and introduce corresponding attacks that expose fundamental flaws in their design.
Analyses
@@
Flaw 2: Token Under-Sampling. Modern VideoLLMs inherit the input token limit of their host LLMs. For example, GPT-4 allows at most 8,192 tokens per input (achiam2023gpt4). Since this token budget is shared between visual tokens and textual tokens, VideoLLMs must compress the tokens of each frame to meet the token budget limit. Formally, given a video of $N$ sampled frames and per-frame token number $P$, the total token number should satisfy:
@@
where $B$ denotes the LLM’s input token limit. Therefore, the token number is reduced to $P^{\prime}$ ($P^{\prime}<P$) through token compression. Many recent works adopt simple downsampling techniques, such as bilinear interpolation (zhang2025videollama3; li2024llava-onevision) or average pooling (zhang2024LLaVA-Video), which aggregates visual tokens on a 2D spatial token grid. For example, an original $14\times 14$ token grid is downsampled to $7\times 7$, retaining only 25% of the spatial tokens.
@@
Flaw 3: Modality Fusion Imbalance. After projection into the language model’s embedding space, visual tokens are often underutilized during decoding. As a result, the LLM tends to prioritize textual information while downplaying or even ignoring signals from the visual encoder, preventing harmful cues from being reflected in the final response. Prior studies (fu2025hidden) show that the standalone visual encoder surpasses the fully fused image-text model on vision-centric benchmarks, underscoring a structural imbalance where visual information loses influence after fusion and barely shapes the final representation. The problem persists in existing VideoLLMs, which reuse image-based encoders to generate visual tokens. Even when these encoders flag harmful content in the sampled frames, their signals are diminished during decoding, hindering the model from faithfully reporting it.
To validate this, we conduct a comparative experiment using LLaVA-Video-7B-Qwen2 and its underlying visual encoder, SigLIP. Following (fu2025hidden), we examine the effectiveness of the visual encoder through a visual probing strategy. Specifically, we construct evaluation videos by inserting harmful content into benign source videos (details in the next section) and randomly sample 100 benign and 100 harmful examples for each of three categories: violence, crime, and pornography. For each video, we examine the binary classification accuracy of both the standalone visual encoder and the full VideoLLM under identical inputs. Figure 3 reports the proportion of correctly identified videos for each category. For benign videos, both the visual encoder and the full VideoLLM achieve comparably high detection rates. However, for the three harmful categories, the full VideoLLM exhibits a significant performance drop compared to the visual encoder. This discrepancy provides concrete empirical support for this flaw, confirming that modality fusion does suppress visual signals even when they are preserved at the visual encoder level.
@@
Threat Model
We assume a strict zero-query black-box setting: the adversary has no knowledge of the targeting VideoLLM’s internals, such as architecture, weights, training data, temporal sampling rate, token compression strategy, or modality fusion scheme, and cannot repeatedly query the model for optimization. The only prior is the knowledge of the three architectural flaws identified previously. The adversary may insert self-sourced harmful clips, but these must remain visible to a human i.e., not single-frame flashes or imperceptible perturbations, so that any detection failure reflects a true omission. This zero-query, training-free attacking setup enables efficient, real-time deployment without per-video adaptation.
Three Attacks
@@
Frame-Replacement Attack (FRA $\rightarrow$ Flaws 1, 3). This attack replaces a segment of the original video with a harmful video clip at a randomly chosen position. Specifically, we select a random insertion point and overwrite the subsequent $t_{r}$ seconds ($t_{r}>1$) with a preselected harmful video clip. Since VideoLLMs employ the sparse uniform sampling, a short replacement window $t_{r}$ is seldom sampled. For instance, in a 2-minute video at 30 FPS, i.e., 3,600 frames, taking only 16 evenly spaced frames gives a stride of 8 seconds (240 frames). A 4-second harmful clip can therefore fit entirely between two sampled frames, leaving the model with no evidence of it, even though human viewers see this harmful segment clearly.
Picture-in-Picture Attack (PPA $\rightarrow$ Flaws 2, 3). We particularly embed a harmful clip in a fixed Picture-in-Picture (PiP) region within each frame, e.g., the bottom-right corner, while the rest of the frame remains unchanged. The PiP region occupies $\eta H\times\eta W$ pixels of each frame, where $\eta\in(0,1)$ is chosen not too small to ensure the malicious content is visible to humans. Because VideoLLMs compress the tokens of each frame to fit the token budget, small peripheral regions are often discarded, failing to influence the model’s output despite being clearly visible.
@@
Table 1: Attack performance. Metric: HOR (%).
Experiments
Experimental Setup
Video Samples. We randomly sample 200 original videos from the LLaVA-Video-178K dataset (zhang2024LLaVA-Video), a widely used benchmark for evaluating VideoLLMs. For harmful clips, we focus on three representative categories: violence, crime, and pornography. These categories are commonly encountered in safety-critical scenarios and easily recognizable to humans. Harmful videos are collected from public datasets (RLVS (soliman2019violence), XD-Violence (wu2020xd-violence), Pornography dataset (avila2013pornography)) and online platforms including YouTube and Pornhub. For each category, we randomly select 10 harmful clips for the three attacks. For each attack, we randomly pair every original video with one harmful video from the corresponding category.
@@
Experimental Results
Attack Effectiveness. Table 1 demonstrates the effectiveness of the three zero-query black-box attacks. For each attack, we fix a reference hyperparameter setting that ensures the harmful content remains clearly visible to humans while inducing substantial omissions by VideoLLMs. Attackers can readily tune these parameters to suit their own objectives.
@@
Visualizations. Figure 5 illustrates representative video examples for all three attacks and harmful content categories. In every case, the injected clip is clearly visible to humans, yet every evaluated model fails to mention it. These omissions reveal the vulnerability of current VideoLLMs to harmful content injection, driven by their fundamental weaknesses.
@@
Discussion
Potential Mitigations. Several directions may help mitigate harmful content omission in VideoLLMs caused by design flaws. Improving frame sampling, for instance, through relevance-based selection (cheng2024focuschat), can slightly increase the chance of capturing harmful segments. Another approach is to perform auxiliary image-level checks using pretrained MLLMs. Since these models typically do not employ token compression, they are better at detecting fine-grained harmful signals, although this substantially increases computational cost. Finally, increasing visual weight during modality fusion may also enhance sensitivity to visual cues. We test denser sampling, relevance-based sampling, and VLM-assisted detection, which offer limited mitigation, with HOR remaining as high as 71% – 95%. This is because coarse sampling before detection is unavoidable (processing all frames is computationally infeasible), allowing harmful frames to be overlooked.
Long Video Understanding. Recent advances have introduced VideoLLMs designed for long video understanding (tens of minutes to hours) (wang2025seal; zhang2025flash). However, the key design flaws identified in this paper still persist. For example, these models continue to rely on sparse temporal sampling, which leaves them vulnerable to the same omission issues observed in shorter videos. As shown in our analysis of harmful clip duration in Appendix, the minimum duration required for a harmful segment to evade all sampled frames increases with video length. This scaling effect makes it even easier to insert undetected harmful content in long-form videos, thereby posing greater security risks. Given the high deployment cost and the current immaturity of long VideoLLMs, we leave a detailed investigation of their safety properties to future work.
Proprietary Models. Although our study focuses on open-source VideoLLMs, the revealed design flaws may still persist in proprietary MLLMs (openaigpt4o) due to similar preprocessing and architectures. For example, Gemini 1.5-Pro also adopts uniform sampling of 16 frames per video (team2023gemini). A thorough investigation of harmful content omission in proprietary models is left for future work.
Other Prompts. We experiment with more informative prompts, such as “Describe any violent scenes.”, but models still fail to detect the harmful content. In FRA, for instance, the failure stems from the fact that harmful frames may be never sampled, making any prompt ineffective. Moreover, even when models respond affirmatively, follow-up questions about the time or location of the harmful content often yield incorrect answers, suggesting that the actual omission rate may exceed what our HOR metric captures.
Conclusion
This work identifies and systematically analyzes three fundamental design flaws in current VideoLLMs: sparse uniform sampling, which leaves large portions of the video unchecked; token under-sampling, which leads to the loss of localized spatial information; and modality fusion imbalance, which suppresses visual signals even when harmful content is captured by the encoder. To demonstrate the consequences of these flaws, we propose three zero-query, black-box attacks that insert harmful content through frame replacement, picture-in-picture, and transparent overlays. Despite the content being readily noticeable to human viewers, these attacks consistently achieve high Harmfulness Omission Rates across multiple mainstream VideoLLMs. This work serves as an early step toward understanding the structural vulnerabilities of VideoLLMs in open-world, safety-critical scenarios. We call for rethinking core design choices and building models that are not only accurate but also safe and reliable.
Acknowledgments
We thank the reviewers for their constructive comments.
Appendix
Overview of Mainstream VideoLLMs
Table 2 summarizes the strategies of frame selection and token compression in mainstream VideoLLMs. As aligned with our core analysis in the main text, most existing models employ uniform frame sampling, either by selecting a fixed number of frames (e.g., 16 or 32) or by applying a low and fixed sampling rate. This design significantly facilitates adversarial attempts to insert harmful content, as large portions of the video are systematically ignored. While a few models incorporate additional frame selection mechanisms, such as keyframe selection (chen2024sharegpt4video) or KNN-based clustering (jin2024chat-univi), the total number of selected frames remains small. Thus, our core conclusion regarding the vulnerability of sparse sampling remains unaffected.
Year Model Frame Selection Token Compression
2024 LLaMA-VID (li2024llama-vid) 1 FPS pooling, 2 tokens per frame
2024 LLaVA-Hound (zhang2024llava-hound) 10 not mentioned
2024 VideoLLaVA (lin2024video-llava) 8 not mentioned
2024 VideoLLaMA 2 (cheng2024videollama2) 8/16 3D Conv & 3D Pool
2024 LLaVA-OneVision (li2024llava-onevision) 32 bilinear interpolation
2024 LLaVA-Video (zhang2024LLaVA-Video) 1 FPS average pooling
2024 LLaVA-NeXT-Video (zhang2024llavanextvideo) 4/8/16/32/64 bilinear interpolation
2024 ShareGPT4Video (chen2024sharegpt4video) 16 not mentioned
2024 Chat-UniVi (jin2024chat-univi) KNN average pooling
2025 Apollo (zohar2025apollo) 2 FPS 16 tokens per frame
2025 NVILA (liu2025nvila) 256 $\rightarrow$ 8 pooling
2025 ViLaMP (cheng2025vilamp) 1 FPS differential feature merging
2025 VideoLLaMA 3 (zhang2025videollama3) 1 FPS bilinear interpolation
2024 LongVLM (weng2024longvlm) 100 hierarchical token merging
2024 VideoStreaming (qian2024videostreaming) 16 average pooling
2025 AKS (tang2025adaptive) 1 FPS not mentioned
2025 FRAG (huang2025frag) 256 $\rightarrow$ 32 not mentioned
Table 2: Overview of frame selection and token compression strategies in mainstream VideoLLMs. FPS: frames per second.
Hyperparameter Analyses
As shown in Figure 7, the probability of missing the inserted segment is high when either the sampling frame count is low or the replaced segment is short. Our results show that when only 4 or 8 frames are sampled, the inserted content is always captured by at most one frame, regardless of duration (from 1% to 10%). In fact, for the inserted segment to appear in more than one sampled frame, its duration must exceed the sampling interval. This becomes particularly critical for VideoLLMs that sample 16 frames, a configuration commonly adopted and often regarded as optimal for minute-long videos (cheng2025vilamp). Under this setting, any segment shorter than 6% of the video duration is guaranteed to be sampled by at most one frame. For a one-minute video, a 4-second harmful segment is sufficient to meet that requirement. Also, such a segment is perceptually salient to human viewers, but is often omitted by VideoLLMs. The analysis above also justifies our experimental setting of $t_{r}=4$ in the main evaluation.
Figure 8: Attack performance under different PiP scaling ratios.